OpenStudy (embracetheweebness98):

Help needed, i'm stuck. What is the simplified form of the quantity x squared minus 5x plus 6 over 15 x y squared all over the quantity 2x squared minus 7x plus 3 over 5 x squared y?

jimthompson5910 (jim_thompson5910):

Sounds like you're dealing with this? \[\Large \frac{x^2-5x+6}{15xy}\div\frac{2x^2-7x+3}{5x^2y}\] Anyways, I'll let @Lord_Box answer this one

OpenStudy (lord_box):

\[\frac{ x^2-5x+6 }{ 15xy^2} \div \frac{ 2x^2-7x+3 }{ 5x^2y }\] Just with a square on the first y.

OpenStudy (lord_box):

Alright when dividing a fraction by another fraction, how is this accomplished?

OpenStudy (embracetheweebness98):

Cross multiplication?

OpenStudy (lord_box):

Yes, multiplying the top of one by the bottom of the other, and vice versa. Can you show me what that would look like?

jimthompson5910 (jim_thompson5910):

Cross multiplication only happens if you set two fractions equal to each other

OpenStudy (embracetheweebness98):

6x2-5x+y+6?

OpenStudy (lord_box):

Alright, I'll clear it up. When you divide one fraction by another, you "flip the second fraction and multiply.\[\frac{ x^2-5x+6 }{ 15xy^2 } \times \frac{ 5x^2y }{ 2x^2-7x+3 }\]

OpenStudy (embracetheweebness98):

okay

jimthompson5910 (jim_thompson5910):

hint: try to factor `x^2-5x+6` and `2x^2-7x+3`

OpenStudy (embracetheweebness98):

the first one is (x-2)(x-3) I couldn't figure out the second

OpenStudy (lord_box):

You can factor the second one by grouping. First, multiply the coefficients of the first and last terms, 2 * 3 = 6. What two numbers multiply to 6 but add to - 7?

OpenStudy (embracetheweebness98):

-6 and -1

OpenStudy (lord_box):

Right, so you can rewrite this as 2x^2 - 6x - x +3. Think of these as two parts, 2x^2 - 6x, then -x + 3, can you identify the common factor that goes into both of these?

OpenStudy (lord_box):

I'll help you out, -x + 3 fits into both of these. Once you find this greatest common factor, you need to take it out of both:\[2x^2-6x-x+3 = 2x(x-3) -1(x-3) = (2x-1)(x-3)\]

OpenStudy (lord_box):

the beginning should read x-3 instead of -x+3, sorry

OpenStudy (embracetheweebness98):

okay

OpenStudy (lord_box):

so right now we have \[\frac{ (x-2)(x-3) }{ 15xy^2 } \times \frac{ 5x^2y }{ (2x-1)(x-3) }\] What all could you cancel out?

OpenStudy (embracetheweebness98):

(x-3)

jimthompson5910 (jim_thompson5910):

see the attached text document to see another way to factor `2x^2-7x+3`

OpenStudy (lord_box):

\[\frac{ (x-2) }{ 15xy^2 } \times \frac{ 5x^2y }{ (2x-1) }\] Good, you cancelled the (x-3), but there's still more that can be simplified. HINT: top right, bottom left

OpenStudy (embracetheweebness98):

Thank you @jim_thompson5910 Oh, okay XD

OpenStudy (embracetheweebness98):

one is 5x^2y and the other is 5xy^2

jimthompson5910 (jim_thompson5910):

Hint: \[\Large \frac{5x^2y}{15xy^2}=\frac{5*x*x*y}{3*5*x*y*y}\] \[\Large \frac{5x^2y}{15xy^2}=\frac{{\color{red}{5}}*{\color{blue}{x}}*x*{\color{purple}{y}}}{3*{\color{red}{5}}*{\color{blue}{x}}*{\color{purple}{y}}*y}\]

OpenStudy (embracetheweebness98):

oh

jimthompson5910 (jim_thompson5910):

That's a detailed way to show that the '5's will pair up and cancel out, a pair of 'x's will cancel out. So will a pair of 'y's. What is left over when all that canceling is done?

OpenStudy (embracetheweebness98):

5x and 3y

jimthompson5910 (jim_thompson5910):

the '5's will cancel

OpenStudy (embracetheweebness98):

oh

jimthompson5910 (jim_thompson5910):

\[\Large \frac{5x^2y}{15xy^2}=\frac{5*x*x*y}{3*5*x*y*y}\] \[\Large \frac{5x^2y}{15xy^2}=\frac{{\color{red}{5}}*{\color{blue}{x}}*x*{\color{purple}{y}}}{3*{\color{red}{5}}*{\color{blue}{x}}*{\color{purple}{y}}*y}\] \[\Large \frac{5x^2y}{15xy^2}=\frac{{\color{red}{\cancel{5}}}*{\color{blue}{x}}*x*{\color{purple}{y}}}{3*{\color{red}{\cancel{5}}}*{\color{blue}{x}}*{\color{purple}{y}}*y}\]

OpenStudy (embracetheweebness98):

yeah, i got confused XD

OpenStudy (embracetheweebness98):

I now have x and 3y

jimthompson5910 (jim_thompson5910):

Yes, so \[\Large \frac{5x^2y}{15xy^2} = \frac{x}{3y}\] which means \[\Large \frac{ (x-2) }{ 15xy^2 } \times \frac{ 5x^2y }{ (2x-1) }\] turns into \[\Large \frac{ (x-2) }{ 3y } \times \frac{ x }{ (2x-1) }\]

OpenStudy (lord_box):

Good stuff. Let's recap: \[\frac{ (x-2) }{ 3y } \times \frac{ x }{ (2x-1) }\]

OpenStudy (lord_box):

\[\frac{ x(x-2) }{ 3y(2x-1)}\]

OpenStudy (embracetheweebness98):

okay, thank you for the help!